OpenStudy (anonymous):

solve each equation. show work please.its a fraction just dnt know how to type on here c/-4 -8 = -42

jhonyy9 (jhonyy9):

(c/-4 )-8= -42 you want writing this ?

OpenStudy (anonymous):

yea

jhonyy9 (jhonyy9):

c+32= 168 c=168-32 c=136

jhonyy9 (jhonyy9):

this is all

jhonyy9 (jhonyy9):

do you understand it ?

jhonyy9 (jhonyy9):

if do you think that this is right what you like to get it please push the good answer button